Circular No. 9032

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 F4 (McNAUGHT) R. H. McNaught reports his discovery of a strongly condensed comet

on CCD images taken with the 0.5-m Uppsala Schmidt reflector at Siding Spring (discovery observation tabulated below), the object showing some diffuseness. Stacked follow-up images by McNaught on Mar. 20.6–20.8 UT show the comet to appear softer than the images of nearby stars of similar brightness. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, G. Sostero, E. Guido, V. Gonano, and P. Camilleri report that nineteen co-added unfiltered CCD 300-s exposures taken on Mar. 20.1 with a 0.45-m f/4.4 reflector at Remanzacco, Italy, show the object to be slightly diffuse when compared to nearby stars of similar brightness, with a short extension nearly 12′′ long toward the west. R. Ligustri, Talmassons, Udine, Italy, reports that eleven stacked 180-s images taken remotely with a 0.25m f/3 reflector near Mayhill, NM, U.S.A., on Mar. 20.4 shows the object’s images to have a FHWM that is 30 percent larger than stars of similar brightness.

2009 UT α_{2000} δ_{2000} Mag. Observer
Mar. 19.68311 15^{h}47^{m}03.^{s}25 −23^{o}10′41′′.8 17.8 McNaught

The available astrometry, the following very preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-F54.

T = 2012 Jan. 10.967 TT  ω = 256.^{o}890
Ω = 53.687 }2000.0
q = 5.80600 AU  i = 80.138

COMETS C/2008 X13, C/2008 X14, C/2008 Y4, C/2008 Y5 (SOHO)

Further to IAUC 9031, additional Kreutz sungrazers have been found on SOHO website images. K. Battams notes that C/2008 X13 was quite diffuse (mag ∼ 7.5) with a very short, diffuse tail. C/2008 X14 was extremely faint (mag ∼ 8.5) and slightly diffuse. C/2008 Y4 and C/2008 Y5 were tiny and stellar in appearance (mag ∼ 8–8.5).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 X13 Dec. 10.729 17^{h}14.^{m}0 −24^{o}52′ C2 RK 2009-E61
C/2008 X14  13.663 17 28.5 −25 04 C2 ZX 2009-E63
C/2008 Y4  19.712 17 56.4 −25 08 C2 BZ 2009-E63
C/2008 Y5  19.896 17 57.7 −25 08 C2 ZJ 2009-E63
2009 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green